### 08 Dec 2023 v3

# Graph neural networks with polynomial activations have limited expressivity

Abstract.
The expressivity of Graph Neural Networks (GNNs) can be entirely characterized by appropriate fragments of the first order logic. Namely, any query of the two variable fragment of graded modal logic (GC2) interpreted over labeled graphs can be expressed using a GNN whose size depends only on the depth of the query. As pointed out by [${}^{+}$20, Gro21], this description holds for a family of activation functions, leaving the possibility for a hierarchy of logics expressible by GNNs depending on the chosen activation function. In this article, we show that such hierarchy indeed exists by proving that not all GC2 queries can be expressed by GNNs with polynomial activation functions. This implies a separation between polynomial and popular non-polynomial activations (such as Rectified Linear Units (ReLUs)), and answers an open question formulated by [Gro21].
1. Introduction
Graph Neural Networks (GNNs) are deep learning architectures for input data that incorporates some relational structure represented as a graph, and have proven to be very performant and efficient for various types of learning problems [ZAL18, ${}^{+}$17, ${}^{+}$19, DBV16, ${}^{+}$08, Ham20, ${}^{+}$15, ${}^{+}$17, ${}^{+}$20, ${}^{+}$20, ${}^{+}$16, ${}^{+}$21, ${}^{+}$20]. Understanding their expressive power, and its dependence on the activation function of the underlying neural networks is one of the basic tasks towards a rigorous study of their computational capabilities.
The notion of expressivity may bear several interpretations based on different mathematical definitions. One of them consists in comparing GNNs to other standard computation models on graphs, such as the color refinement or the Wesfeiler-Lehman algorithms. Namely, it has been proven [${}^{+}$19, XHLJ18] that the color refinement algorithm precisely captures the expressiveness of GNNs in the following sense: there is a GNN distinguishing two nodes of a graph if and only if color refinement assigns different colors to these nodes. This description of the expressivity of GNNs may not be uniform, since it allows the size of the underlying neural networks to grow with the size of the input graphs. Furthermore, if the activation functions are supposed to be piecewise polynomial, this property has been proved to be not uniform [KB23]: in order to express color refinement, the size of the GNNs has to grow with the size of the input graphs. In a similar fashion, explicit lower bounds on the size of the underlying neural network(s) of a GNN expressing the color refinement can be derived for piecewise-polynomial activation functions. This can be derived from the construction presented in [KB23] and upper bounds on the number of regions of a neural network with a piecewise polynomial activation function.
The second angle of attack to understand the expressivity of GNNs is to describe the types of boolean queries that GNNs characterize. More precisely, given a boolean query interpreted over the space of graphs, does there exist a GNN whose output characterizes the output of this boolean query? For example, can a GNN express if a node of a graph is part of a clique of given size? From that perspective, the expressive power of GNNs can be surprisingly exactly characterized by a suitable formal logic, called the two variable fragment of graded model logic (GC2). Any GNN expresses a query of this logic, and conversely, any query of this logic can be expressed by a GNN. Moreover, this characterization is now uniform: the size of a GNN required to express a GC2 query does not depend on the size of the input graph, but depends only on the number of subformulas (or depth) of the query. This is in contrast with the aforementioned notion of expressivity, comparing GNNs to color refinement.
This discrepancy can be explained at a high level as follows. On the one hand, the first notion of expressivity is global: it compares the expressive power of GNNs to a general powerful algorithm whose capacity to distinguish graphs is close to graph isomorphism testing [CFI92]. On the other hand, the second notion of expressivity is a local one: it considers “local” queries that are considerably weaker than the color refinement algorithm to distinguish graphs or vertices. The known proofs for the characterization using the GC2 logic [${}^{+}$20, Gro21] provide an explicit construction of a GNN expressing a given query. However, this constructive approach does not work for any activation function, leading to the following question: does there exist a hierarchy of logics expressible by GNNs depending on their activation function?
In this article, we present a partial answer to this question: we prove that GNNs with any polynomial activations cannot express all GC2 queries, although GNNs with popular activations functions that are not polynomial can express any GC2 query, a result known since [${}^{+}$20]. This answers an open question formulated by [Gro21] in his survey about the logic of GNNs. Interestingly, our result also aligns with the limitations of polynomial activations for approximating continuous functions on compact sets with neural networks [LLPS93].
The rest of this article is organized as follows. In Section 2, we present the definitions needed to formally state our results in Section 3. The proofs of technical lemmata and our main result is given in Section 4.
2. Preliminaries
In order to state what is meant by a query and the two variable fragment of graded modal logic, we first need to collect some basic definitions in formal logic (Section 2.1). We then present the fragments of logics we are interested for the study of the expressivity of GNNs. In Section 2.2 we provide the formal definition of a GNN as well as the notion of expressing a query for a GNN. For an excellent introduction to descriptive complexity theory from the angle of modern graph theory, we refer the reader to [Gro17].
We assume graphs to be finite, undirected, simple, and vertex-labeled: a graph is a tuple $G=(V(G),E(G),P_{1}(G),...,P_{\ell}(G))$ consisting of a finite vertex set $V(G)$, a binary edge relation $E(G)\subset V(G)^{2}$ that is symmetric and irreflexive, and unary relations $P_{1}(G),\cdots,P_{\ell}(G)\subset V(G)$ representing $\ell>0$ vertex labels. In the following, the number $\ell$ of labels, which we will also call colors, is supposed to be fixed and does not grow with the size of the input graphs. When there is no ambiguity about which graph $G$ is being considered, $N(v)$ refers to the set of neighbors of $v$ in $G$ not including $v$. $\lvert G\rvert$ will denote the number of vertices of $G$.
2.1. Finite variable counting logic
Definition 1.
A first order logic is given by a countable set of symbols, called the alphabet of the logic:
  1. (1)
    Boolean connectives: ¬,,,$\neg$,$\vee$,$\wedge$,$\rightarrow$,$\leftrightarrow$ , ∨ , ∧ , → , ↔
  2. (2)
    Quantifiers: $\forall$,$\exists$ , ∃
  3. (3)
    Equivalence/equality symbol: $\equiv$
  4. (4)
    Variables: $x_{0},x_{1},\ldots$ (finite or countably infinite set)
  5. (5)
    Punctuation: ((((, )))) and ,.
  6. (6)
    1. (a)
      A (possibly empty) set of constant symbols.
    2. (b)
      For every natural number $n\geq 1$, a (possibly empty) set of $n$-ary function symbols.
    3. (c)
      For every natural number $n\geq 1$, a (possibly empty) set of $n$-ary relation symbols.
Remark 1.
Items 1-5 are common to any first order logic. Item 6 changes from one system of logic to another. Example: In Graph theory, the first order logic has:
  • no constant symbols
  • no function symbol
  • a single $2$-ary relation symbol $E$ (which is interpreted as the edge relation between vertices). When graphs are supposed labeled with $\ell$ colors: $\ell$ function symbols $\mathsf{col}_{1},\cdots,\mathsf{col}_{\ell}$. $\mathsf{col}_{i}(v\in G)$ returns $1$ if the color of $v$ is the $i$-th one.
The set of symbols from Item 6 is called the vocabulary of the logic. It will be denoted by $S$ and the first order logic based on $S$ will be denoted by $FO(S)$.
Definition 2.
The set of terms in a given first order logic $FO(S)$ is a set of strings over the alphabet defined inductively as follows:
  1. (1)
    Every variable and constant symbol is a term.
  2. (2)
    If $f$ is an $n$-ary function symbol, and $t_{1},\ldots,t_{n}$ are terms, then $f(t_{1},\ldots,t_{n})$ is a term.
Definition 3.
The set of formulas in a given first order logic is a set of strings over the alphabet defined inductively as follows:
  1. (1)
    If $t_{1},t_{2}$ are terms, then $t_{1}\equiv t_{2}$ is a formula.
  2. (2)
    If $R$ is an $n$-ary relation symbol, and $t_{1},\ldots,t_{n}$ are terms, then $R(t_{1},\ldots,t_{n})$ is a formula.
  3. (3)
    If $\phi$ is a formula, then $\neg\phi$ is a formula.
  4. (4)
    If $\phi_{1},\phi_{2}$ are formulas, then $(\phi_{1}\vee\phi_{2})$, $\phi_{1}\wedge\phi_{2}$, $\phi_{1}\rightarrow\phi_{2}$ and $\phi_{1}\leftrightarrow\phi_{2}$ are formulas.
  5. (5)
    If $\phi$ is a formula and $x$ is a variable, then $\forall x\phi$ and $\exists x\phi$ are formulas.
The set of all variable symbols that appear in a term $t$ will be denoted by $\operatorname{var}(t)$. The set of free variables in a formula is defined using the inductive nature of formulas:
  1. (1)
    $\operatorname{free}(t_{1}\equiv t_{2})=\operatorname{var}(t_{1})\cup\operatorname{var}(t_{2})$
  2. (2)
    $\operatorname{free}(R(t_{1},\ldots,t_{n}))=\operatorname{var}(t_{1})\cup\ldots\cup\operatorname{var}(t_{n})$
  3. (3)
    $\operatorname{free}(\neg\phi)=\operatorname{free}(\phi)$
  4. (4)
[…]
  5. (5)
    $\operatorname{free}(\forall x\phi)=\operatorname{free}(\phi)\setminus\{x\}$
  6. (6)
    $\operatorname{free}(\exists x\phi)=\operatorname{free}(\phi)\setminus\{x\}$
Remark 2.
The same variable symbol may be a free symbol in $\phi$, but appear bound to a quantifier in a subformula of $\phi$.
Definition 4.
The set of sentences in a first order logic are all the formulas with no free variables, i.e., $\{\phi:\operatorname{free}(\phi)=\emptyset\}$.
Definition 5.
Given a first order logic $FO(S)$, an $S$-structure is a pair $\mathcal{U}=(A,I)$ where $A$ is a nonempty set, called the domain/universe of the structure, and $I$ is a map defined on $S$ such that
  1. (1)
    $I(c)$ is an element of $A$ for every constant symbol $c$.
  2. (2)
    $I(f)$ is a function from $A^{n}$ to $A$ for every $n$-ary function symbol $f$.
  3. (3)
    $I(R)$ is a function from $A^{n}$ to $\{0,1\}$ (or equivalently, a subset of $A^{n}$) for every $n$-ary relation symbol $R$.
Given an $S$-structure $\mathcal{U}=(A,I)$ for $FO(S)$, an assignment is a map from the set of variables in the logic to the domain $A$. An interpretation of $FO(S)$ is a pair $\mathcal{U},\beta$, where $\mathcal{U}$ is an $S$-structure and $\beta$ is an assignment.
We say that an interpretation $\mathcal{U},\beta$ satisfies a formula $\phi$, if this assignment restricted to the free variables in $\phi$ evaluates to 1, using the standard Boolean interpretations of the symbols of the first order logic in Items 1-5 above.
Example 1.
Consider the first order logic of graphs where $S=\{E\}$ where $E$ is a $2$-ary relation symbol. The following formula expresses that no vertex is isolated: $\forall x\exists yE(x,y).$ Similarly, the formula $\forall x\neg E(x,x)$ expresses the fact that we do not want any self loops. A more interesting example is
$\psi(x):=\forall x\exists y\exists z(\neg(y\equiv z)\wedge E(x,y)\wedge E(x,z))$
which expresses that the graph must be undirected. We can augment that formula as follows. Let $\phi$ be defined as
$\phi(x):=\forall x\exists y\exists z(\neg(y\equiv z)\wedge E(x,y)\wedge E(x,z))\wedge\forall w(E(x,w)\rightarrow((w\equiv y)\vee(w\equiv z)))$
$\phi$ expresses that a node $x$ of a graph has exactly to out-neigbhors.
Hence, the formula:
$\forall x\forall y(E(x,y)\leftrightarrow E(y,x))$ $\quad\wedge\quad\psi(x)$
expresses that the graph must be undirected and every vertex has exactly two neighbors.
Any (directed or undirected) graph $G=(V,E)$ gives an $S$-structure for the above logic: $A=V$ and we have the $2$-ary edge relation $x,y\in A$ are related if and only if $(x,y)\in E$. An assignment is a map from the set of variables in the logic to $V$.
Any formula in $FO(S)$ can be thought of as a $0/1$ function on the class of all interpretations of $FO(S)$:
  1. (1)
    If $\phi$ is a sentence, then any graph $G$ is an $S$-structure and is mapped to 00 or $1$, depending on whether $G$ verifies $\phi$ or not.
  2. (2)
    If $\phi(x)$ is a formula with a single free variable $x$, then any pair $(G,v)$, where $G=(V,E)$ is a graph and $v\in V$, is an interpretation with $G$ as the $S$-structure and the assignment $\beta$ maps $x$ to $v$. Thus, every pair $(G,v)$ is mapped to 00 or $1$, depending on whether $(G,\beta)$ verifies $\phi$ or not. This example can be extended to handle formulas with multiple free variables, where we may want to model $0/1$ functions on subsets of vertices.
Remark 3.
We now see that sentences in $FO(S)$ can be thought of as $0/1$ functions on any class of $S$-structures, and thus they can be thought of as classifiers on that class of $S$-structures. Similarly, formulas with free variables also model $0/1$ functions on pairs of $S$-structures and elements or tuples from the domain of the structure.
Definition 6.
[…]
This is equivalent to the depth of the tree representing the formula, based on the inductive definition. The length/size of the formula is the total number nodes in this tree. Up to constants, this is the number of leaves in the tree, which are called the atoms of the formula.
Definition 7 (Guarded model logic GC [${}^{+}$20, Gro21]).
The fragment of guarded logic GC is formed using quantifiers that restrict to range over the neighbours of the current nodes. GC-formulas are formed from the atomic formulas by the Boolean connectives and quantification restricted to formulas of the form $\exists^{\geq p}(E(x,y)\land\psi))$, where x and y are distinct variables and $x$ appears in $\psi$ as a free variable. Note that every formula of GC has at least one free variable. We are interested in the 2-variable fragment GC2, also known as graded modal logic.
A more constructive definition is given as follows: a graded modal logic formula is either $\mathsf{Col}(x)$ (returning 1 or 0 for one of the palette colors) or one of the following:
$\lnot\phi(x),\quad\phi(x)\land\psi(x),or\quad\exists^{\geq N}y(E(x,y)\land\phi(y))$
where $\phi$ and $\psi$ are GC2 formulas and $N$ is a positive integer:
Example 2 ([${}^{+}$20]).
[…]
$\gamma$ queries if $x$ has blue color and as at least one neigbhor which has two blue neighbors. Then $\gamma$ is in GC2. Now,
[…]
is not in GC2 because the use of the guard $\lnot E(x,y)$ is not allowed. However,
[…]
is in GC2 because the negation $\lnot$ is applied to a formula in GC2.
2.2. GNNs and expressing queries
Definition 8 (Graph Neural Network (GNN)).
A GNN is a recursive embedding of vertices of a labeled graph by relying on the underlying adjacency information and node features. Each vertex $v$ is attributed an indicator vector $\xi^{0}(v)$ of size $\ell$, encoding the color of the node $v$: the colors being indexed by the palette $\{1,\cdots,\ell\}$, $\xi^{0}(v)=e_{i}$ (the $i$-th canonical vector) if the color of the vertex $v$ is $i$. The GNN is fully characterized by:
$\circ$ A combination function $\mathsf{comb}:\mathbb{R}^{2\ell}\longrightarrow\mathbb{R}^{\ell}$ which is a feedforward neural network with given activation function $\sigma:\mathbb{R}\longrightarrow\mathbb{R}$.
$\circ$ The update rule of the GNN at iteration $t\in\mathbb{N}$ for any labeled graph $G$ and vertex $v\in V(G)$, is given as follows:
$\xi^{0}(v)\text{ is the indicator vector of the color of $v$},\qquad\xi^{t+1}(v)=\mathsf{comb}(\xi^{t}(v),\sum_{w\in N(v)}\xi^{t}(w))$
We say that a GNN has polynomial activations provided the underlying neural network $\mathsf{comb}$ has polynomial activation functions.
Remark 4.
A first addition to the basic model of GNNs in Definition 8 is global readout, to account for information at a global level, in the update rule:
$\xi^{t+1}(v):=\mathsf{comb}(\xi^{t}(v),\sum_{w\in N(v)}\xi^{t}(w),\sum_{w\in V(G)}\xi^{t}(w))$
A second variant of GNNs consists in adding the possibility for various $\mathsf{comb}$ functions, depending on the iteration $t$. Our results also extend to such GNNs, but for ease of presentation, we restrict to this definition. See also Remark 5 for more details on the logic of GNNs with global readouts.
Definition 9 ([Gro21]).
Suppose that $\xi$ is the vertex invariant computed by a GNN. We say that a GNN expresses a unary query if the GNN expresses the query $Q$ if there is an $\epsilon<\frac{1}{2}$ such that for all graphs G and vertices $v\in V(G)$.
$\begin{cases}\xi(G,v)\geq 1-\epsilon&\text{if $v\in Q(G)$}\\ \xi(G,v)\leq\epsilon&\text{if $v\notin Q(G)$}\end{cases}$
3. Formal statement of the results
We are now equipped to state the known results regarding the expressivity of GNNs:
Theorem 1.
[${}^{+}$20, Gro21] Let Q be a unary query expressible in graded modal logic GC2. Then there is a GNN whose size depends only on the depth of the query, that expresses Q.
Theorem 1 has a partial converse, with a slight weakening:
Theorem 2.
[${}^{+}$20] Let Q be a unary query expressible by a GNN and also expressible in first-order logic. Then Q is expressible in GC2.
Remark 5.
The right logic for GNNs with global readout is not GC2, but the extension of the first order logic with counting quantifiers $\exists^{\geq p}$, denoted by C. Let C2 be the fragment of C consisting of all formulas with at most 2 variables. Then, we have the following result: [${}^{+}$20]: Let Q be a Boolean or unary query expressible in C2. Then there is a GNN with global readout that expresses Q.
In contrast with Theorem 1, we prove that GNN with polynomial activations cannot express any GC2 query:
Theorem 3.
There are GC2 queries that cannot be expressed by any GNN (with or without global readout) with polynomial activations.
4. Proofs
To prove our result we explicitely construct a query that no GNN with polynomial activation (and whose size does not depend on the input graphs) can express. We can then easily extend the reasoning to more queries. We prove this statement by contradiction: on the one hand we evaluate a GNN with polynomial activations on a set of given inputs; and on the other hand, interprete our query on the same set of input graphs. We show that the polynomial obtained by the first evaluation cannot verify the constraints imposed by the query.
Our set of inputs are formed using rooted unicolored trees of the form shown in Figure 1 which is a tree of depth two whose depth one vertices have prescribed degrees $k_{1},\cdots,k_{m}$, with $k_{1},\cdots,k_{m}\geq 1$. We first collect three elementary Lemmata, one that will be useful to extract monomials of largest degree in a multivariate polynomial (Lemma 1). Since the trees are parameterized by $m$-tuples of integers $k_{1},\ldots,k_{m}$, the embedding of the root node computed by the GNN at any iteration is a function of these $m$ integers. Since the activations are polynomials, these embeddings of the root node are multivariate symmetric polynomial functions of $k_{1},\ldots,k_{m}$ (Lemma 2). Lemma 3 state the degree boundedness such polynomial obtained by a GNN on these tree structures. Our proof of Theorem 3 builds on these results combined with fundamental properties of symmetric multivariate polynomials of bounded degree.
[…]
Figure 1. $T[k_{1},\cdots,k_{m}]$
Lemma 1.
Let $p$ be a positive integer and let $S\subset\mathbb{N}^{p}$ be a finite subset of integral vectors of the nonnegative orthant, such that $\lvert S\rvert\geq 2$. Then there exist $x^{*}\in S$ and $u\in\mathbb{N}^{p}$ such that for any $x\in S-\{x^{*}\}$, $\langle x^{*},u\rangle>\langle x,u\rangle$.
Proof.
Consider one vector $x^{*}$ maximizing $\lVert x\rVert_{2}^{2}$ over $S$, i.e. $\lVert x^{*}\rVert_{2}^{2}=\mathsf{max}_{x\in S}\lVert x\rVert_{2}^{2}$ then let $x\in S-\{x^{*}\}$. Such $x^{*}$ and $x$ exist because $S$ is finite and $\lvert S\rvert\geq 2$.
- Case 1: $x$ is not colinear to $x^{*}$. It follows from the Cauchy-Schwarz inequality, that $\langle x^{*},x\rangle<\lVert x^{*}\rVert_{2}\lVert x\rVert_{2}$. Hence
$\langle x^{*},x^{*}-x\rangle$ $=\lVert x^{*}\rVert_{2}^{2}-\langle x^{*},x\rangle>\lVert x^{*}\rVert_{2}(\lVert x^{*}\rVert_{2}-\lVert x\rVert_{2})>0$
- Case 2: $x\in S-\{x^{*}\}$ is colinear to $x^{*}$, i.e. $x=\lambda x^{*}$ with $\lambda\in\mathbb{R}$. Since $x^{*}$ is maximizing the $2$-norm on S, then $10\leq\lambda<$ 1. Then
$\langle x^{*},x^{*}-x\rangle=\lVert x^{*}\rVert(1-\lambda)>0$
In both cases, $\langle x^{*},x^{*}-x\rangle>0$. Hence, we can set $u:=x^{*}\in S\subset\mathbb{N}^{p}$, and the Lemma is proved. ∎
Lemma 2.
Let $\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ be the embedding of the tree displayed in Figure 1 obtained via a GNN with polynomial activation functions after $t$ iterations, where $\xi^{0}(v)=1$ for all vertices $v\in V(T[k_{1},\ldots,k_{m}])$. Then, for any iteration $t$, there exists a symmetric multivariate polynomial $F$ such that $\xi^{t}(T[k_{1},\ldots,k_{m}],s)=F(k_{1},\cdots,k_{m})$. Furthermore, the degree of $F$ does not depend on $m$, but only on the underlying neural network and $t$.
Proof.
We first prove by induction on $t$ that, for any vertex $v\in V(T[k_{1},\cdots,k_{m}])$, $\xi^{t}(T[k_{1},\ldots,k_{m}],v)$ is a polynomial function of the $k_{i}$’s.
Base case: for $t=0$ this is trivial since all vertices are initialised with the constant polynomial 1, whose degree does not depend on $m$.
Induction step: Suppose the property is true at iteration $t$, i.e for each node $w$, $\xi^{t}(T[k_{1},\ldots,k_{m}],w)$ is a multivariate polynomial of the $k_{i}$’s. Since
$\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)=\phi(\xi^{t}(T[k_{1},\ldots,k_{m}],v),\sum_{w\in N(v)}\xi^{t}(T[k_{1},\ldots,k_{m}],w))$
where $\phi$ is a multivariate polynomial of $k_{1},\cdots,k_{m}$.
By composition $\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)$ is a multivariate polynomial of $k_{1},\cdots,k_{m}$. By induction the degree of $\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)$ depends only on $t$ and the degree of $\phi$, which does not depend on $m$.
Finally, we know from [Gro21, Theorem VIII.1] that the color refinement algorithm refines any GNN at any iteration. Since the tuple obtained by color refinement for the vertex $s$ is invariant with respect to permutations of the $k_{i}$’s, $\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ is also invariant with respect to permutations of the $k_{i}$’s. ∎
Lemma 3.
Let $\xi^{t}(T[k_{1},\cdots,k_{m}],v)$ be the embedding of a vertex $v$ of the tree $T[k_{1},\cdots,k_{m}]$ displayed in Figure 1, returned by a GNN with polynomial activations after $t$ iterations. Then, for any vertex $v\in V(T[k_{1},\cdots,k_{m}])$, the degree of the sequence of polynomials $P_{m}:=\xi^{t}(T[k_{1},\cdots,k_{m}],v)$ is bounded, i.e. for any iteration $t$, there exists a positive integer $q_{t}$ (independent of $m$ and $v$) such that for any positive integer $m$, $\mathsf{deg}(P_{m})\leq q_{t}$.
Proof.
We will prove the statement by induction on the number of iterations.
Base case: At the first iteration ($t=0$), $P_{m}$ is constant equal to $1$ ($q_{1}=1$) for any $m$.
Induction step: Suppose that for any iteration $t\leq T$ , there exists $q_{t}\in\mathbb{N}$ (that does not depend on $m$ nor the vertex $v\in V(T[k_{1},\cdots,k_{m}])$), such that for any integer $m$, and for any iteration $t$, $\mathsf{deg}(P_{m})\leq q_{t}$. Then, using the update rule:
$\underbrace{\xi^{t+1}(T[k_{1},\cdots,k_{m}],v)}_{Q_{m}}=\phi(\underbrace{\xi^{t}(T[k_{1},\cdots,k_{m}],v)}_{R_{m}},\underbrace{\sum_{w\in N(v)}{}\xi^{t}(T[k_{1},\cdots,k_{m}],w)}_{S_{m}})$
$R_{m}$ and $S_{m}$ are multivariate polynomials of $m$ variables. By the induction hypothesis, there exist $r_{t}$ and $s_{t}$, such that for any $m$, $\mathsf{deg}(R_{m})\leq q_{t}$ and $\mathsf{deg}(S_{m})\leq q_{t}$.
$\phi$ is a bivariate polynomial of degree independent of $m$ (neural network with a polynomial activation function). Let $v$ be its degree. Hence, the degree of $Q_{m}$ is at most $v\times q_{t}^{2}$. Hence the property remains true at $t+1$, setting $q_{t+1}:=v\times q_{t}^{2}$. ∎
Proof of Theorem 3.
Consider the following query of GC2:
(1) $Q(s)=\lnot\left(\exists^{\geq 1}x(E(s,x)\land\exists^{\leq 1}sE(x,s))\right)=\forall xE(s,x)\exists^{\geq 2}sE(x,s)$
$Q$ is true if and only if all the neigbhors of the node $s$ have degree at least $2$. We will prove by contradiction that any bounded GNN with a polynomial activation function cannot express the query $Q$. Let $P_{m}:=\xi^{t}(T[k_{1},\cdots,k_{m}],s)$ be the embedding of the source node of $T[k_{1},\cdots,k_{m}]$ returned by a GNN with polynomial activations, after a fixed number of iterations $t$. Suppose that it can express the query $Q$, then;
$\begin{cases}P_{m}(k_{1},\cdots,k_{m})\geq 1-\epsilon&\text{if $s\in Q(T[k_{1},\cdots,k_{m}])$}\\ P_{m}(k_{1},\cdots,k_{m})\leq\epsilon&\text{if $s\notin Q(T[k_{1},\cdots,k_{m}])$}\end{cases}$
Let $\tilde{P}_{m}:=P_{m}-\frac{1}{2}$ and $\epsilon^{\prime}:=\frac{1}{2}-\epsilon$. Interpreting the query $Q$ over $T[k_{1},\cdots,k_{m}]$ implies the following constraints on the sequence of polynomial $\tilde{P}_{m}$:
(2) $\exists\epsilon^{\prime}>0\;\text{ such that:}\;\begin{cases}\forall k\in\mathbb{N}^{m},\exists i\in\{1,\cdots,m\},k_{i}=0\implies\tilde{P}_{m}(k_{1},\cdots,k_{m})\leq-\epsilon^{\prime}&\\ \forall k\in\mathbb{N}^{m},\forall i\in\{1,\cdots,m\},k_{i}>0\implies\tilde{P}_{m}(k_{1},\cdots,k_{m})\geq\epsilon^{\prime}&\end{cases}$
Let $q$ be the degree of $P_{m}$, $q:=\mathsf{deg}(P_{m})=\mathsf{deg}(\tilde{P}_{m})$ and let $S$ be the set of exponents of the monomials of $P_{m}$, i.e.
$S:=\{(\alpha_{1},\cdots,\alpha_{m})\in\mathbb{N}^{m}:\alpha_{1}+\cdots+\alpha_{m}\leq q\text{ and $k_{1}^{\alpha_{1}}\cdots k_{m}^{\alpha_{m}}$ is a monomial of $\tilde{P}_{m}$}\}$
First, note that $q\geq 2$ ($\tilde{P}_{m}$ cannot be linear, otherwise its zero locus would contain a union of several hyperplanes as soon as $m\geq 2$). This insures that $\lvert S\rvert\geq 2$.
Then using Lemma 1 (with $p=m$) tells us there exists $\alpha^{*}\in S$ and $u=(u_{1},\cdots,u_{m})\in\mathbb{N}^{m}$ such that for any $\alpha^{\prime}\in S-\{\alpha^{*}\}$,
$\langle\alpha^{*},u\rangle>\langle\alpha^{\prime},u\rangle$
Claim 1: The (univariate) monomial $t^{\langle\alpha^{*},u\rangle}$ is the monomial of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m}})$ of largest degree.
Proof.
$\tilde{P}_{m}=\sum_{\alpha\in S}\gamma_{\alpha}k_{1}^{\alpha_{1}}\cdots k_{m}^{\alpha_{m}}\implies$ $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},t^{u_{m}})=\sum_{\alpha\in S}\gamma_{\alpha}t^{\langle\alpha,u\rangle}$
Hence, the monomial of largest degree of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m}})$ is the one such that $\langle\alpha,u\rangle$ is (strictly) maximized when $\alpha\in S$. By construction, it is $\alpha^{*}$. ∎
Now, suppose that $\tilde{P}_{m}$ has bounded degree (w.r.t $m$). Namely, any monomial of largest degree of $\tilde{P}_{m}$ does not contain all the variables. Without loss of generality (by symmetry of $\tilde{P}_{m}$), we can suppose that the monomial $k_{1}^{\alpha^{*}_{1}}\cdots k_{m}^{\alpha^{*}_{m}}$ does not contain $k_{m}$, i.e. $\alpha^{*}_{m}=0$.
Claim 2: In these conditions, the (univariate) monomial $t^{\langle\alpha^{*},u\rangle}$ is also the monomial of of largest degree of $P_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},0)$.
Proof.
Evaluating $\tilde{P}_{m}$ in $(t^{u_{1}},\cdots,t^{u_{m-1}},0)$ removes the contribution of each monomial of $P_{m}$ containing the last variable, and keeps only the contribution of the monomials containing it:
$\tilde{P}_{m}=\sum_{\alpha\in S}\gamma_{\alpha}k_{1}^{\alpha_{1}}\cdots k_{m}^{\alpha_{m}}\implies$ $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},0)=\sum_{\begin{subarray}{c}\alpha\in S\\ \alpha=(\alpha_{1},\cdots,\alpha_{m-1},0)\end{subarray}}\gamma_{\alpha}t^{\langle\alpha,u\rangle}$
Therefore, the monomial of largest degree of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},0)$ is the one such that $\langle\alpha,u\rangle$ is (strictly) maximized when $\alpha\in S$ and $\alpha_{m}=0$. By construction, such $\alpha$ is $\alpha^{*}$. ∎
However, Conditions 2 imply that the leading monomial of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m}})$ must have a positive coefficient but the leading monomial of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},0)$ must have a negative coefficient, by taking both limits when $t$ tends to +$\infty$+ ∞ (this limit can be constant). This is a contradiction because Claims 1 and 2 imply that these coefficients are both $\gamma_{\alpha^{*}}$.
Therefore, if $P_{m}$ expresses $Q$, then the monomials of greatest degree of $\tilde{P}_{m}$ contain all the variables, hence $\mathsf{deg}(P_{m})=\mathsf{deg}(\tilde{P}_{m})\geq m$. This gives a contradiction with Lemma 3 since $P_{m}=\tilde{P}_{m}+\frac{1}{2}$ is supposed to be computed by a GNN. Therefore, no GNN with polynomial activations can express the query $Q$.
Remark 6.
Note that the proof can easily be extended to a larger family of queries. Namely, for any integer $p\geq 2$, let
$Q_{p}(s):=\lnot\left(\exists^{\geq 1}x(E(s,x)\land\exists^{\leq(p-1)}sE(x,s))\right)=\forall xE(s,x)\exists^{\geq p}sE(x,s)$
$Q_{p}$ queries if vertex $s$ has neighbors whose degree are all at least $p$. Then any $(Q_{p})_{p\in\mathbb{N}}$ cannot be expressed by any GNN with polynomial activations. We leave the question of determining precisely the queries that can be expressed with these polynomial activations for future work.
Remark 7.
The proof of Theorem $3$ was initially attempted using queries of the form:
$\tilde{Q}_{p}(s)=\lnot\left(\exists^{\geq 1}x(E(s,x)\land\exists^{\geq(p+1)}sE(x,s))\right)=\forall xE(s,x)\exists^{\leq p}sE(x,s)$
Which expresses that all the neighbors of $s$ have degree at most $k$. Note the similarity between $Q_{p}$ from Remark 6 and $\tilde{Q}_{p}$. Although $\tilde{Q}_{p}$ also seems a good candidate that cannot be expressed by a GNN with polynomial activations, we could not conclude with the same approach as in the proof of Theorem 3, due to the following interesting fact:
There exists $\epsilon>0$ and a sequence of symmetric polynomial $(p_{m})_{m\in\mathbb{N}}\in\mathbb{R}[x_{1},\cdots,x_{m}]$ of bounded degree (i.e. there exists an integer $q$ such that for any $m$, $\mathsf{deg}(p_{m})\leq q$) and for any $m,p_{m}$ is greater than $\epsilon$ on the vertices of the unit hypercube $\{0,1\}^{m}$, and less than $-\epsilon$ on all the other points of $\mathbb{N}^{m}$. $p_{m}=1-\sum_{i=1}^{m}x_{i}^{2}+\sum_{i=1}^{m}x_{i}^{4}$ is an example of such sequence of symmetric polynomials.
5. Acknowledgements
The author thanks Amitabh Basu, Edinah K. Gnang, and Floris Geerts for insightful discussions. The author gratefully acknowledges support from Air Force Office of Scientific Research (AFOSR) grant FA95502010341.