### 31 Jan 2024 v5

# Graph neural networks with polynomial activations have limited expressivity

Abstract.
The expressivity of Graph Neural Networks (GNNs) can be entirely characterized by appropriate fragments of the first order logic. Namely, any query of the two variable fragment of graded modal logic (GC2) interpreted over labeled graphs can be expressed using a GNN whose size depends only on the depth of the query. As pointed out by [${}^{+}$20, Gro21], this description holds for a family of activation functions, leaving the possibibility for a hierarchy of logics expressible by GNNs depending on the chosen activation function. In this article, we show that such hierarchy indeed exists by proving that GC2 queries cannot be expressed by GNNs with polynomial activation functions. This implies a separation between polynomial and popular non polynomial activations (such as Rectified Linear Units) and answers an open question formulated by [Gro21].
1. Introduction
Graph Neural Networks (GNNs) are deep learning architectures for input data that incorporates some relational structure represented as a graph, and have proven to be very performant and efficient for various types of learning problems [ZAL18, ${}^{+}$17, ${}^{+}$19, DBV16, ${}^{+}$08, Ham20, ${}^{+}$15, ${}^{+}$17, ${}^{+}$20, ${}^{+}$20, ${}^{+}$16, ${}^{+}$21, ${}^{+}$20]. Understanding the expressive power of GNNs, and its dependence on the activation function of the underlying neural networks is one of the basic tasks towards a rigorous study of their computational capabilities.
In this context, several approaches have been conducted in order to describe and characterize the expressivity of GNNs. The first approach consists in comparing GNNs to other standard computation models on graphs such as the color refinement or the Wesfeiler-Leman algorithms. This reduction type of approach stands to reason as the computational models of GNNs, Wesfeiler-Leman/color refinement algorithms are intimately connected: they all fall under the paradigm of trying to discern something about the global structure of a graph from local neighborhood computations. In that regard, it has been proven [${}^{+}$19, XHLJ18] that the color refinement algorithm precisely captures the expressivity of GNNs. More precisely, there is a GNN distinguishing two nodes of a graph if and only if colour refinement assigns different colours to these nodes. This results holds if one supposes that the size of the underlying neural networks are allowed to grow with the size of the input graph. Hence, in his survey, [Gro21] emphasizes the fact that this equivalence has been established only for unbounded GNN, and asks: Can GNNs with bounded size simulate color refinement? In [${}^{+}$22], the authors answer by the negative if the underlying neural network are supposed to have Rectified Linear Unit (ReLU) activation functions. In [KB23] the authors provide a generalization of this result, for GNNs with piecewise polynomial activation functions. Furthermore, explicit lower bounds on the neural network size to simulate the color refinement can be derived for piecewise-polynomial activation functions given upper bounds on the number of regions of a neural network with piecewise-polynomial activation.
The second line of research to study the expressive power of GNNs is to characterize the types of boolean queries interpreted over labeled graphs that a GNN can simulate. Given a boolean query $Q$ (taking as input a graph, or a graph and one of its vertices), does there exist a GNN and whose output characterizes the output of $Q$? For example, can a GNN express if a vertex of a graph is part of a clique of given size? Furthermore, can we characterize the set of queries that can be simulated by GNNs? In that context if the size of the GNN depends only on the to grow with the size of the input, graph, then we say that the GNN expresses the query uniformly. Uniform expressivity is interessing from a practical standpoint as it captures the expressivity of GNNs of fixed size with respect to the input graphs.
Several mathematical answers to such questions have been already obtained. Notably, a complete description of the logic of the queries that the GNNs can express uniformly have been proved. This suitable formal logic interpreted over labelled graphs is a two variable fragment of graded model logic (GC2). Any GNN expresses a query of this logic, and conversely, any query of this logic can be expressed by a GNN whose size and iterations only depends on the depth of the query [${}^{+}$20, Gro21]. For specific activation functions such as ReLUs, the size of a GNN required to express a given query of GC2 does not depend on the size of the input graph, but only on the number of subformulas (or depth) of the query. The known proofs of this result [${}^{+}$20, Gro21] provide an explicit construction of a GNN with ReLU activations that expresses a given query. In recent results [Gro23], the author provides a more general description of the logics expressible by GNNs, and also treat the non-uniform case. The uniform case is obtained for rational piecewise-linear activations (or equivalently, rational ReLUs), and a non-uniform result is presented for GNNs with general arbitrary real weights and activation functions. The author also presents new results about the expressivity of GNNs if random initialisation is allowed on the features of the vertices. In this article, we focus on uniform expressivity and consider the following question: What is the impact of the activation and aggregation functions on the logic uniformly expressed by GNNs?
Main contribution. In this article we show a separation between polynomial and non-polynomial activations (and in particular, piecewise linear activations) with respect to the logic expressible by GNNs with those activation functions. More precisely, we prove that GNNs with polynomial activation and aggregation functions cannot express all GC2 queries (uniformly), although GNNs with piecewise linear activations and a linear aggregation function can. This results holds even if: i) the weights of the polynomial GNNs are arbitrary real numbers with infinite precision, and ii) the weights of the GNNs with piecewise polynomial are restricted to integers (also, the underlying neural networks are supposed to have finitely many linear pieces). This shows how the power of graph neural networks can change immensely if one changes the activation function of the neural networks. Our result constitutes an additional step towards an complete understanding of the impact of the activation function on the formal expressivity of GNNs.
The rest of this article is organized as follows. Section 2 presents the definitions of GNNs and the background logic. In Section 3, we state our main result and compare it to the existing ones. Section 5 presents an overview of the proof of our main result, as well as proofs of the technical lemmata are presented. We conclude with some remarks and open questions in Section 6.
2. Preliminaries
2.1. Graph Neural Networks (GNNs)
We assume the input graphs of GNNs to be finite, undirected, simple, and vertex-labeled: a graph is a tuple $G=(V(G),\allowbreak E(G),P_{1}(G),\allowbreak\cdots,P_{\ell}(G))$ consisting of a finite vertex set $V(G)$, a binary edge relation $E(G)\subset V(G)^{2}$ that is symmetric and irreflexive, and unary relations $P_{1}(G),\allowbreak\cdots,P_{\ell}(G)\subset V(G)$ representing $\ell>0$ vertex labels. In the following, we suppose that the $P_{i}(G)$’s form a partition of the set of vertices of $G$, i.e. each vertex has a unique label. Also, the number $\ell$ of labels, which we will also call colors, is supposed to be fixed and does not grow with the size of the input graphs. This setup introduced by [${}^{+}$20, Gro21] allow to model the presence of features of the vertices of input graphs in practical applications. In order to describe the logic of GNNs, we also take into account access to the color of the vertices into the definition of the logic considered, as we shall see in Section 2.2. When there is no ambiguity about which graph $G$ is being considered, $N(v)$ refers to the set of neighbors of $v$ in $G$ not including $v$. $\lvert G\rvert$ will denote the number of vertices of $G$. We use simple curly brackets for a set $X=\{x\in X\}$ and double curly brackets for a multiset $Y=\{\{y\in Y\}\}$.
Definition 1 (Neural network).
Fix an activation function $\sigma:\mathbb{R}\to\mathbb{R}$. For any number of hidden layers $k\in\mathbb{N}$, input and output dimensions $w_{0}$, $w_{k+1}\in\mathbb{N}$, a $\mathbb{R}^{w_{0}}\rightarrow\mathbb{R}^{w_{k+1}}$ neural network with $\sigma$ activation is given by specifying a sequence of $k$ natural numbers $w_{1},w_{2},\cdots,w_{k}$ representing widths of the hidden layers and a set of $k+1$ affine transformations $T_{i}:\mathbb{R}^{w_{i-1}}\rightarrow\mathbb{R}^{w_{i}}$, $i=1,\ldots,k+1$. Such a NN is called a $(k+1)$-layer NN, and is said to have $k$ hidden layers. The function $f:\mathbb{R}^{w_{0}}\rightarrow\mathbb{R}^{w_{k+1}}$ computed or represented by this NN is:
$f=T_{k+1}\circ\sigma\circ T_{k}\circ\cdots T_{2}\circ\sigma\circ T_{1}.$
In the following, the Rectified Linear Unit activation function $\operatorname{ReLU}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is defined as $\operatorname{ReLU}(x)=\max(0,x)$. The Sigmoid activation function $\operatorname{Sigmoid}:\mathbb{R}\rightarrow(0,1)$ is defined as $\operatorname{Sigmoid}(x)=\frac{1}{1+e^{-x}}.$
Definition 2 (Graph Neural Network (GNN)).
A GNN is a recursive embedding of vertices of a labeled graph by relying on the underlying adjacency information and node features using a neural network (cf. Definition 1). Each vertex $v$ is attributed an indicator vector $\xi^{0}(v)$ of size $\ell$, encoding the color of the node $v$: the colors being indexed by the palette $\{1,\cdots,\ell\}$, $\xi^{0}(v)=e_{i}$ (the $i$-th canonical vector) if the color of the vertex $v$ is $i$. A GNN is characterized by:
  • $\circ$
    A positive integer $T$ for the number of iterations of the GNN (also called the number of layers of the GNN), positive integers $(d_{t})_{t\in\{1,\allowbreak\cdots,T\}}$ and $(d^{\prime}_{t})_{t\in\{0,\cdots,T\}}$ for inner dimensions. $d_{0}=d^{\prime}_{0}=\ell$ is the input dimension of the GNN (number of colors) and $d_{T}$ is the output dimension.
  • $\circ$
    a sequence of combination and aggregation functions $(\mathsf{comb}_{t},\allowbreak\mathsf{agg}_{t})_{t\in\{1,\cdots,T\}}$. Each aggregation function $\mathsf{agg}_{t}$ maps each finite multiset of vectors of $\mathbb{R}^{d_{t-1}}$ to a vector in $\mathbb{R}^{d^{\prime}_{t}}$. For any $t\in\{1,\cdots,T\}$, each combination function $\mathsf{comb}_{t}:\mathbb{R}^{d_{t-1}+d^{\prime}_{t}}\longrightarrow\mathbb{R}^{d_{t}}$ is a neural network with given activation function $\sigma:\mathbb{R}\longrightarrow\mathbb{R}$.
The update rule of the GNN at iteration $t\in\{0,\cdots,T-1\}$ for any labeled graph $G$ and vertex $v\in V(G)$, is given by:
$\xi^{t+1}(v)=\mathsf{comb}(\xi^{t}(v),\mathsf{agg}\{\{\xi^{t}(w))\,:\,w\in\mathcal{N}_{G}(v)\}\})$
We say that a GNN has polynomial activations provided the underlying neural network $\mathsf{comb}$ has polynomial activation functions.
Remark 1.
The type of GNN in our Definition 2 is sometimes referred to as aggregation-combine GNNs without global readout. Here are a few variants that can be found in the litterature:
  • Recurrent GNNs, allowing $\mathsf{comb}_{t}$ and $\mathsf{agg}_{t}$ functions, do not depend on the iteration $t$.
  • GNNs with global readout, which allow the aggregation functions to also take as input the embeddings of all the vertices of the graph. See Remark 4 for more details on the logic side of things.
  • General Message-passing GNNs that allow operations before the aggregation on the neigbhors as well as the current vertex. As mentioned in [Gro23][Remark 4.1], it is not clear whether this type of GNN increases expressivity.
2.2. Logical background: GC2 and queries
The first-order language of graph theory we consider is built up in the usual way from a vocabulary containing variables $x_{1},x_{2},\cdots,x_{m}$, the relations symbols $E$ and $=$, the logical connectives ,,¬$\land$,$\lor$,$\lnot$,$\rightarrow$ , ∨ , ¬ , →, and the quantifiers $\forall$ and $\exists$.
Given a number of colors $\ell$ and a colored graph $G=(V(G),E(G),\allowbreak P1(G),...,P^{\ell}(G))$, where $P1(G),\cdots,P^{\ell}(G)\subset V(G)$ represent the vertex colors (one color per vertex),
  • The universe $A$ of the logic is given by $A=V$.
  • the set of symbols $S$ of the first order logic we consider is composed of:
    • $\circ$
      the $2$-ary edge relation symbol $E$: $(x,y)\in A^{2}$ are related if and only if $(x,y)\in E$.
    • $\circ$
      function symbols $\mathsf{col}_{1},\cdots,\mathsf{col}_{\ell}$. $\mathsf{col}_{i}(v\in G)$ returns 1 if the color of $v$ is the $i$-th one.
  • The pair $(A,I)$ allows us to interpret the first order logic, where the map $I$ is a function from $A^{2}$ to $\{0,1\}$ (or equivalently, a subset of the set of pair of vertices).
The quantifiers, variables, and set of symbols form the alphabet of the logic considered. The set of formulas in the logic is a set of strings over the alphabet defined inductively (see the appendix for a formal definition). The pair $(A,I)$, called an $S$-structure for the logic, allows us to interpret the formulas of the logic (here, in the space of graphs). Any graph is naturally associated to an $S$ structure in order to interpret the sentences of formula of the logic. For example, the following formula interpreted over a graph $G$ expresses that no vertex of $G$ is isolated: $\forall x\exists yE(x,y).$ Similarly, the formula $\forall x\neg E(x,x)$ expresses the fact that we do not want any self loops. A more interesting example is given by
(1) $\psi:=\forall x\forall y[E(x,y)\rightarrow E(y,x)\land x\neq y]$
expresses that $G$ is undirected and loop-free. Similarly,
(2) $\begin{split}\phi:=\forall x\exists y\exists z(\neg(y=z)\wedge E(x,y)\wedge E(x,z))\wedge\forall w(E(x,w)\rightarrow((w=y)\vee(w=z)))\end{split}$
expresses that every node $x$ of the considered graph has exactly two out-neigbhors.
Since GNNs can output values for every vertex of a graph, the formulas we will be using to have to take as “input” some vertex variable; we call such variable a free variable. Concretely, a free variable is a variable which is not bound to a quantifier. Namely, the Formulas 1 and 2, contain no free variable. However, the formula $\phi(x):=\exists y\exists z(\neg(y=z)\wedge E(x,y)\wedge E(x,z))$ has a single free variable $x$. We interpret $\phi$ with $(G,v)$ as a $S$-structure, where $G$ is a graph and $v$ one of its vertices (and the assignment that maps $x$ to $v$). In this case, $\phi$ expresses that vertex $v$ has two out-neighbors in $G$.
Any formula in $FO(S)$ (the first order logic on $S$) can be thought of as a $0/1$ function on the class of all interpretations of $FO(S)$:
  1. (1)
    If $\phi$ is a sentence, then any graph $G$ is an $S$-structure and is mapped to 00 or $1$, depending on whether $G$ verifies $\phi$ or not.
  2. (2)
    If $\phi(x)$ is a formula with a single free variable $x$, then any pair $(G,v)$, where $G=(V,E)$ is a graph and $v\in V$, is an interpretation with $G$ as the $S$-structure and the assignment $\beta$ maps $x$ to $v$. Thus, every pair $(G,v)$ is mapped to 00 or $1$, depending on whether $(G,\beta)$ verifies $\phi$ or not. This example can be extended to handle formulas with multiple free variables, where we may want to model $0/1$ functions on subsets of vertices.
[…]
Definition 3 (Guarded model logic (GC) and GC2 [${}^{+}$20, Gro21]).
The fragment of guarded logic GC is formed using quantifiers that restrict to range over the neighbours of the current nodes. GC-formulas are formed from the atomic formulas by the Boolean connectives and quantification restricted to formulas of the form $\exists^{\geq p}y(E(x,y)\land\psi))$, where x and y are distinct variables and $x$ appears in $\psi$ as a free variable. Note that every formula of GC has at least one free variable. We refer to the 2-variable fragment of GC as GC2, also known as graded modal logic (with two variables).
A more constructive definition can be given as follows: a graded modal logic formula is either $\mathsf{Col}(x)$ (returning 1 or 0 for one of the palette colors) or one of the following:
$\lnot\phi(x),\quad\phi(x)\land\psi(x),or\quad\exists^{\geq N}y(E(x,y)\land\phi(y))$
where $\phi$ and $\psi$ are GC2 formulas and $N$ is a positive integer:
Example 1 ([${}^{+}$20]).
[…]
$\gamma$ queries if $x$ has blue color and as at least one neigbhor which has two blue neighbors. Then $\gamma$ is in GC2. Now,
[…]
is not in GC2 because the use of the guard $\lnot E(x,y)$ is not allowed. However,
[…]
is in GC2 because the negation $\lnot$ is applied to a formula in GC2.
2.3. Color refinement
For the proof of our main result, we will need two additional definitions that we include for completeness:
Definition 4 (Embeddings and refinement).
Given a set $X$, an embedding $\xi$ is a function that takes as input a graph $G$ and a vertex $v\in V(G)$, and returns an element $\xi(G,v)\in X$. We say that an embedding $\xi$ refines an embedding $\xi^{\prime}$ if and only if for any graph $G$ and any $v\in V(G),\xi(G,v)=\xi(G,v^{\prime})\implies\xi^{\prime}(G,v)=\xi^{\prime}(G,v^{\prime})$. When the graph $G$ is clear from context, we use $\xi(v)$ as shorthand for $\xi(G,v)$.
Definition 5 (Color refinement).
Given a graph $G$, and $v\in V(G)$, let $(G,v)\mapsto\mathsf{col(G,v)}$ be the function which returns the color of the node $v$. The color refinement refers to a procedure that returns a sequence of embeddings $cr^{t}$, computed recursively as follows:
- $cr^{0}(G,v)=\mathsf{col}(G,v)$
- For $t\geq 0$, $\mathsf{cr}^{t+1}(G,v):=(\mathsf{cr}^{t}(G,v),\{\{\mathsf{cr}^{t}(G,w):w\in N(v)\}\})$
In each round, the algorithm computes a coloring that is finer than the one computed in the previous round, that is, $\mathsf{cr}^{t+1}$ refines $\mathsf{cr}^{t}$. For some $t\leq n:=\lvert G\rvert$, this procedure stabilises: the coloring does not become strictly finer anymore.
The following connection between color refinement and GNNs will be useful to prove our main result. Notably, the theorem holds regardless of the choice of the aggregation function $\mathsf{agg}$ and the combination function $\mathsf{comb}$.
Theorem 1 ([${}^{+}$19, XHLJ18]).
Let $d$ be a positive integer, and let $\xi$ be a vertex invariant computed by an GNN after $d$ iterations. Then $\mathsf{cr}^{d}$ refines $\xi$, that is, for all graphs $G,G^{\prime}$ and vertices $v\in V(G)$, $v^{\prime}\in V(G^{\prime})$, $\mathsf{cr}^{(d)}(G,v)=\mathsf{cr}^{d}(G^{\prime},v^{\prime})=\xi(G,v)=\xi(G^{\prime},v^{\prime})$.
3. Formal statement of the results
Definition 6 ([Gro21]).
Suppose that $\xi$ is the vertex embedding computed by a GNN. We say that a GNN expresses uniformly a unary query $Q$ there is a real $\epsilon<\frac{1}{2}$ such that for all graphs G and vertices $v\in V(G)$.
$\begin{cases}\xi(G,v)\geq 1-\epsilon&\text{if $v\in Q(G)$}\\ \xi(G,v)\leq\epsilon&\text{if $v\notin Q(G)$}\end{cases}$
We are now equipped to state the known previous results regarding the expressivity of GNNs:
Theorem 2.
[${}^{+}$20, Gro21] Let Q be a unary query expressible in graded modal logic GC2. Then there is a GNN whose size depends only on the depth of the query, that expresses Q uniformly.
Remark 2.
$\ell$ being the number of colors of the vertices in the input graphs, the family of GNNs with $\mathsf{agg}=\mathsf{sum}$, and $\mathsf{comb}(x,y)=\operatorname{ReLU}(Ax+By+C)$ (where $A\in\mathbb{N}^{\ell\times\ell}$, $B\in\mathbb{N}^{\ell\times\ell}$ and $C\in\mathbb{N}^{\ell}$) is sufficient to uniformly express all queries of GC2 uniformly. Furthermore, for each query $Q$ of depth $q$, there is a GNN of this type with at most $q$ iterations that expresses uniformly $Q$.
Example 2.
Let $Q$ be the following GC2 query:
[…]
$A=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 1&0&1&0\end{pmatrix},B=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\end{pmatrix},c=\begin{pmatrix}0\\ 0\\ 0\\ -1\end{pmatrix}$
and let $\sigma$ be the clipped ReLU function, i.e. $\sigma(\cdot):=\min(1,\max(0,\cdot))$ (the clipped ReLU can be computed by a neural network with ReLU activations). Then, it can be verified that $Q$ can be computed in $4$ iterations with the update rule:
$\xi^{0}(v)=1,\quad\xi^{t+1}(v):=\sigma(A\xi^{t}(v)+B(\sum_{w\in N(v)}\xi^{t}(w)))$
i.e. $Q_{i}(v)=\xi^{4}(v)_{i}$. In particular, $Q_{i}(v)=1\iff\xi^{4}(v)_{i}=1$. Note that in that case, we are able to compute exactly the query.
Remark 3.
The ability of GNNs to compute exactly GC2 queries is a general situation, as shown in the proof of Theorem 2 included in the first section of the appendix. (We also report that in the constructive proof of [${}^{+}$20], the matrices $A$ and $B$ should be replaced by their transpose). We emphasize here that one cannot mimic the proof for sigmoid activations, even by replacing exact computation by uniform expressivity.
Theorem 2 has a partial converse, with a slight weakening:
Theorem 3.
[${}^{+}$20] Let Q be a unary query expressible by a GNN and also expressible in first-order logic. Then Q is expressible in GC2.
Remark 4.
The right logic for GNNs with global readout (cf. Remark 1) is not GC2. If $C2$ is the fragment of the first order logic with counting quantifiers $(\exists^{\geq p})$ and with at most 2 variables; then we have the following result [${}^{+}$20]: Let Q be a Boolean or unary query expressible in C2. Then there is a GNN with global readout that expresses Q.
In contrast with Theorem 2, we prove:
Theorem 4.
There are GC2 queries that no GNN with polynomial activations and aggregations can uniformly express.
[…]
4. Proof that piecewise linear GNNs (and AGG=SUM) are as expressive as GC2
We present here a proof of Theorem 2 for completeness and as a warmup for our main result.
Proof of Theorem 2.
One can reformulate the claim of the Theorem as follows.
[…]
The overall GNN will take as input the graph $G$ as well as for each node of $v$, $\xi^{0}(v)\in\{0,1\}^{\ell}$ encoding the colors of each node of $G$; and after $L$ iterations, outputs for each node a vector $\xi^{d}(v)\in\{0,1\}^{d}$. Furthemore, at each intermediate iteration $t\in\{1,\cdots,d-1\}$, the constructed GNN will verify $\xi^{t}(v)\in\{0,1\}^{d}$. This property will be crucial for the inductive argument to go through.
In order to prove the claim, we simply need to find appropriate $(\mathsf{comb}_{t})_{1\leq t\leq d}$ and $(\mathsf{agg}_{t})_{1\leq t\leq d}$ functions such that if $\xi^{t}$ verifying the update rule:
$\xi^{t+1}(G,v)=\mathsf{comb}_{t}(\xi^{t}(G,v),\mathsf{agg}_{t}(\{\{\xi^{t}(G,v)):v\in\mathcal{N}_{G}(v)\}\}))$
then $\xi^{t}$ have the property we are seeking, i.e. it computes the given query $Q$. We will prove that we can find such $\mathsf{comb}_{t}$ and $\mathsf{agg}_{t}$ functions by induction on the depth of $Q$.
Base case. If $Q$ has depth $1$, $Q$ is one of $\ell$ color queries, and this can be computed via a GNN in one iteration, whose underlying neural network is the projection onto the $i-$th coordinate, i.e.
- $\mathsf{comb}_{0}(x,y)=\text{proj}_{i}(x)$
- $\mathsf{agg}_{0}$ can be chosen as any aggregation function.
Induction step. Let suppose $Q$ be a query of depth $d>1$.
First construction. By the induction hypothesis, we here suppose that we have access to some $(\mathsf{comb}_{t})_{1\leq t\leq d-1}$ and $(\mathsf{comb}_{t})_{1\leq t\leq d-1}$ such that for any $j\in\{1,\cdots,d-1\}$ and for any $1\leq i\leq j$, $Q_{j}(v)=1\iff\xi_{j}^{i}(v)=1$. Recall that $\text{sub}(Q)=(Q_{1},Q_{2},\cdots,Q_{d})$, i.e. the $Q_{i}$s form an enumeration of the sub-formulas of $Q$. In particular, $Q_{d}=Q$. In the following construction, $\xi^{i}$ keeps “in memory” the output of the subformulas $Q_{j}$, for $j\leq i$. For each case described above , we show that we are able to construct $\mathsf{comb}_{d}$ and $\mathsf{agg}_{d}$ in the following form:
- $\mathsf{comb}_{d}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},(x,y)\mapsto\sigma(A_{d}X+B_{d}Y+c_{d})$, where $\sigma(\cdot)=\min(1,\max(0,\cdot))$ is the clipped ReLU. Note that a clipped ReLU can be computed by a Neural Network with ReLU activations.
- $\mathsf{agg}_{d}$ is the sum function.
Due to the inductive nature of GC2, either one of the following holds:
  • Case 1: there exist subformulas $Q_{j}$ and $Q_{k}$ of $Q$, such that $\ell(Q_{j})+\ell(Q_{k})=d$ and $Q=Q_{j}\land Q_{k}$
  • Case 2: $Q(x)=\lnot Q_{j}(x)$ where $Q_{j}$ is a query of depth $d-1$
  • Case 3: there exists a subformula $Q_{j}$ of $Q$ such that $\ell(Q_{j})=d-1$ and $Q(x)=\exists^{\geq N}y(E(x,y)\land Q_{j}(y))$
We will first give general conditions on the update of the combinations and aggregate functions, and then conclude that these conditions are actually be met by constant $\mathsf{comb}$ and $\mathsf{agg}$ functions:
- $A_{d}$ gets the same first $d-1$ rows as $A_{d-1}$.
- $B_{d}$ gets the same first $d-1$ rows as $B_{d-1}$
- $c_{d}$ get the same first $d-1$ coordinates as $c_{d-1}$,
  • Case 1: The $d$-th row of $A_{d}$ gets all zeros except: $(A_{d})_{jd}=1$, $(A_{d})_{kd}=1$. and the $d$-th row of $B_{t}$. Set $(c_{d})_{d}=0$.
  • Case 2: The $d$-th row of $A_{d}$ gets all zeros except: $(A_{d})_{jd}=1$. The $d$-th row of $B_{d}$ is set to 00 except $(B_{d})_{jd}=-1$. Set $(c_{d})_{d}=-1$.
  • Case 3: The $d$-th row of $A_{d}$ gets all zeros except: $(A_{d})_{jd}=1$. The $d$-th row of $B_{d}$ is set to 00 except $(B_{d})_{jd}=-1$. Set $(c_{d})_{d}=-N+1$.
What remains to prove is the following:
$\text{For any $i\in\{1,\cdots,d\}$},\quad\xi_{i}^{d}(G,v)=1\iff Q_{i}(v)=1$
Due to our update rule described for each case, the first $d-1$ coordinates of $\xi^{d}(G,v)$ are the same as $\xi^{d-1}(G,v)$. Hence, the property is true for $i\leq d-1$ by immediate induction. We are left to show that $\quad\xi_{d}^{d}(G,v)=1\iff Q(v)=Q_{d}(v)=1$. Here, we use the fact that:
  • for every node $v$, every coordinate of $\xi^{d-1}(v)$ is in $\{0,1\}$.
  • $\sigma$ is the clipped ReLU activation function: $\sigma(\cdot)=\min(1,\max(0,\cdot))$.
Since $\xi^{d}(v)=\sigma(A_{t}\xi^{d-1}(v)+B_{t}\sum_{w\in N(v)}\xi^{d-1}(w)+c_{t})$ This follows from an immediate discussion on the three cases described previously, and ends the induction on $d$.
An important feature of the update described above is that $(A_{d},B_{d},c_{d})$ verifies all the conditions imposed for every $(A_{t},B_{t},c_{t})_{1\leq t\leq d}$ to compute all subformulas $Q_{t}$. The updates of $A_{t}$, $B_{t}$ and $c_{t}$ are only made for the $t$-th row and $t$-th entry, and do not depend on the previous columns but only on the query $Q$. Hence, we may as well start from the beginning by setting $(A_{t},B_{t},c_{t})$ to $(A_{d},B_{d},c_{d})$, instead of changing these matrices at every iteration $t$. In these conditions, the combination function $\mathsf{comb}_{t}$, parametrized by $A_{t}$, $B_{t}$ and $c_{t}$ can be defined independently of $t$. The same holds for $\mathsf{agg}_{t}$ as it can be chosen as the sum for any iteration.
Second construction. We refer to the proof of [Gro21] for an approach where the GNN is non-recurrent (each $\mathsf{comb}_{t}$ in that case depends on $t$). However, we insist on the fact that $\xi^{t}$ is a $\{0,1\}$ is a vector so that the proof with a clipped ReLU activation goes through.
The reason for which the constructive proof in [${}^{+}$20] does not extend to other activations (namely, sigmoid) is that for a fixed positive integer $p$, we are looking for a function $f_{p}:\mathbb{R}\rightarrow\mathbb{R}$ such that for some $0<\epsilon<\frac{1}{2}$, for any integer $N>0$ and for any $x_{1},\cdots,x_{N}\in[0,\frac{1}{2}-\epsilon]\cup[\frac{1}{2}+\epsilon,1]$,
$f_{p}(\sum_{i=1}^{N}x_{i})\geq\frac{1}{2}+\epsilon\iff\text{there are at least$p$ $x_{i}$'s such that $x_{i}\geq\frac{1}{2}+\epsilon$}$
Such $f_{p}$ obviously does not exist: on the one hand, it is necessary that $f(x)\geq\frac{1}{2}+\epsilon$ for any $x\geq p$. Furthermore, it is possible to pick $x_{1},\cdots,x_{N}$ such that for any $i$, $x_{i}\in[0,\frac{1}{2}-\epsilon]$ but $x_{1}+\cdots x_{N}\geq p$. This in turn would imply $f_{p}(\sum_{i=1}^{N}x_{i})\geq\frac{1}{2}+\epsilon$ but all $x_{i}$’s are smaller than $\frac{1}{2}$. The main reason why the constructive proof above goes through is because we restrict the embeddings to $\{0\}\cup\{1\}$ using $f_{p}(.)=\text{cReLU}(\cdot-p+1)$ where cReLU is the clipped ReLU.
5. Proof of our main result
Overview. To prove our result we construct a query that no GNN with polynomial activation can uniformly express. We prove this statement by contradiction: on the one hand we interpret the embedding returned by a GNN with polynomial activations on a set of given input graphs, as a polynomial of some parameters of the graph structure. On the other hand, we interpret our query on the same set of input graphs. We show that if GNN were to uniformly express the query, then the polynomial obtained by the first evaluation cannot verify the constraints imposed by the query. Our approach uses fundamental properties of multivariate polynomials and can easily extend to a large family of queries.
Our set of inputs are formed using rooted unicolored trees of the form shown in Figure 1 which is a tree of depth two whose depth one vertices have prescribed degrees $k_{1},\cdots,k_{m}$, with $k_{1},\cdots,k_{m}\geq 1$. We first collect three elementary Lemmata, one that will be useful to extract monomials of largest degree in a multivariate polynomial (Lemma 1). Since the trees are parameterized by $m$-tuples of integers $k_{1},\ldots,k_{m}$, the embedding of the root node computed by the GNN at any iteration is a function of these $m$ integers. Since the activations are polynomials, these embeddings of the root node are multivariate symmetric polynomial functions of $k_{1},\ldots,k_{m}$ (Lemma 2). Furthermore, the degree of these polynomials is bounded by a constant independent of $m$.Our proof of Theorem 4 builds on these results combined with fundamental properties of symmetric multivariate polynomials of bounded degree.
[…]
Figure 1. $T[k_{1},\cdots,k_{m}]$
Lemma 1.
Let $p$ be a positive integer and let $S\subset\mathbb{N}^{p}$ be a finite subset of integral vectors of the nonnegative orthant, such that $\lvert S\rvert\geq 2$. Then there exist $x^{*}\in S$ and $u\in\mathbb{N}^{p}$ such that for any $x\in S-\{x^{*}\}$, $\langle x^{*},u\rangle>\langle x,u\rangle$.
Proof.
Consider one vector $x^{*}$ maximizing $\lVert x\rVert_{2}^{2}$ over $S$, i.e. $\lVert x^{*}\rVert_{2}^{2}=\mathsf{max}_{x\in S}\lVert x\rVert_{2}^{2}$ then let $x\in S-\{x^{*}\}$. Such $x^{*}$ and $x$ exist because $S$ is finite and $\lvert S\rvert\geq 2$.
- Case 1: $x$ is not colinear to $x^{*}$. It follows from the Cauchy-Schwarz inequality, that $\langle x^{*},x\rangle<\lVert x^{*}\rVert_{2}\lVert x\rVert_{2}$. Hence
$\langle x^{*},x^{*}-x\rangle=\lVert x^{*}\rVert_{2}^{2}-\langle x^{*},x\rangle>\lVert x^{*}\rVert_{2}(\lVert x^{*}\rVert_{2}-\lVert x\rVert_{2})>0$
- Case 2: $x\in S-\{x^{*}\}$ is colinear to $x^{*}$, i.e. $x=\lambda x^{*}$ with $\lambda\in\mathbb{R}$. Since $x^{*}$ is maximizing the $2$-norm on S, then $10\leq\lambda<$ 1. Then
$\langle x^{*},x^{*}-x\rangle=\lVert x^{*}\rVert(1-\lambda)>0$
In both cases, $\langle x^{*},x^{*}-x\rangle>0$. Hence, we can set $u:=x^{*}\in S\subset\mathbb{N}^{p}$, and the Lemma is proved. ∎
Lemma 2.
Let $\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ be the embedding of the tree displayed in Figure 1 obtained via a GNN with polynomial activation and aggregation functions after $t$ iterations, where $\xi^{0}(v)=1$ for all vertices $v\in V(T[k_{1},\ldots,k_{m}])$. Then, for any iteration $t$, there exists a symmetric multivariate polynomial $F$ such that $\xi^{t}(T[k_{1},\ldots,k_{m}],s)=F(k_{1},\cdots,k_{m})$. Furthermore, the degree of $F$ does not depend on $m$, but only on the underlying neural network and $t$.
Proof.
For clarity, we will perform two separate inductions, one for the existence of the symmetric polynomial, and the second for the degree boundedness.
We first prove by induction on $t$ that, for any vertex $v\in V(T[k_{1},\allowbreak\cdots,k_{m}])$, $\xi^{t}(T[k_{1},\ldots,\allowbreak k_{m}],v)$ is a polynomial function of the $k_{i}$’s.
Base case: for $t=0$ this is trivial since all vertices are initialised with the constant polynomial 1, whose degree does not depend on $m$.
Induction step: Suppose the property is true at iteration $t$, i.e for each node $w$, $\xi^{t}(T[k_{1},\ldots,k_{m}],w)$ is a multivariate polynomial of the $k_{i}$’s. Since
$\xi^{t+1}(T[k_{1},\ldots,k_{m}],v)=\mathsf{comb}($ $\xi^{t}(T[k_{1},\ldots,k_{m}],v),$
$\mathsf{agg}(\{\{\xi^{t}(T[k_{1},\cdots,k_{m}],w):w\in N(v)\}\}))$
where $\mathsf{comb}$ is a neural network with polynomial activations, hence a multivariate polynomial. Also, $\mathsf{agg}$ is supposed polynomial in the entries of its multiset argument. Then by composition, $\xi^{t+1}(T[k_{1},\allowbreak\ldots,k_{m}],v)$ is a multivariate polynomial of $k_{1},\cdots,k_{m}$.
To conclude on the symmetry of the polynomial, we need the intermediary claim:
Claim: Let $T[k_{1},\cdots,k_{m}]$ be two rooted trees given by Figure 1, and let $s$ be its source vertex. Then for any iteration $t$, the color refinement embedding $\mathsf{cr}^{t}(s)$ is invariant by permutation of the $k_{i}$’s.
Proof of claim.
By induction of $t\geq 0$, we prove the following claim: for any integer $t\geq 0$, $\mathsf{cr}^{t}(s)$ and the multiset $\{\{\mathsf{cr}^{t}(x_{1}),\allowbreak\cdots,\mathsf{cr}^{t}(x_{m})\}\}$ are invariant by permutation of the $k_{i}$’s.
Base case: $(t=0)$ is obvious since $\mathsf{cr}^{0}(v)=1$ for any vertex $v\in T[k_{1},\cdots,k_{m}]$.
Induction step: Suppose that for some iteration $t$, $\mathsf{cr}^{t}(s)$ and $\{\{\mathsf{cr}^{t}{x_{1}},\cdots,\mathsf{cr}^{t}(x_{m})\}\}$ are invariant by permutation of the $k_{i}$’s. $\mathsf{cr}^{t+1}(s)$ being invariant by permutation of the $k_{i}$’s is a consequence of: $\mathsf{cr}^{t+1}(s)=(\mathsf{cr}^{t}(s),\{\{\mathsf{cr}^{t}(x_{1}),\cdots,\mathsf{cr}^{t}(x_{m})\}\})$
We know from Theorem 1 that the color refinement algorithm refines any GNN at any iteration. Since the tuple obtained by color refinement for the vertex $s$ is invariant with respect to permutations of the $k_{i}$’s, $\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ is also invariant with respect to permutations of the $k_{i}$’s.
Finally, the degree boundedness also follows by induction:
Base case: At the first iteration ($t=0$), $P_{m}$ is constant equal to $1$ ($q_{1}=1$) for any $m$.
Induction step: Suppose that for any iteration $t\leq T$ , there exists $q_{t}\in\mathbb{N}$ (that does not depend on $m$ nor the vertex $v\in V(T[k_{1},\cdots,k_{m}])$), such that for any integer $m$, and for any iteration $t$, $\mathsf{deg}(P_{m})\leq q_{t}$. Then, using again the update rule:
$\underbrace{\xi^{t+1}(T[k_{1},\cdots,k_{m}],v)}_{Q_{m}}=\mathsf{comb}($ $\underbrace{\xi^{t}(T[k_{1},\cdots,k_{m}],v)}_{R_{m}},$
$\underbrace{\mathsf{agg}(\{\{\xi^{t}(T[k_{1},\cdots,k_{m}],w):w\in N(v)\}\})}_{S_{m}})$
$R_{m}$ and $S_{m}$ are multivariate polynomials of $m$ variables. By the induction hypothesis, there exist $r_{t}$ and $s_{t}$, such that for any $m$, $\mathsf{deg}(R_{m})\leq q_{t}$ and $\mathsf{deg}(S_{m})\leq q_{t}$.
$\mathsf{comb}$ is a bivariate polynomial of degree independent of $m$ (neural network with a polynomial activation function). Let $v$ be its degree. Hence, the degree of $Q_{m}$ is at most $v\times q_{t}^{2}$. Hence the property remains true at $t+1$, setting $q_{t+1}:=v\times q_{t}^{2}$. ∎
Proof of Theorem 4.
Consider the following query of GC2:
$Q(s)=\lnot\left(\exists^{\geq 1}x(E(s,x)\land\exists^{\leq 1}sE(x,s))\right)=\forall xE(s,x)\exists^{\geq 2}sE(x,s)$
$Q$ is true if and only if all the neigbhors of the node $s$ have degree at least $2$. We will prove by contradiction that any bounded GNN with polynomial activations and aggregations cannot uniformly express the query $Q$. Let $P_{m}:=\xi^{t}(T[k_{1},\cdots,k_{m}],s)$ be the embedding of the source node of $T[k_{1},\cdots,k_{m}]$ returned by a GNN with polynomial activations, after a fixed number of iterations $t$. Suppose that it can uniformly express the query $Q$, then;
$\begin{cases}P_{m}(k_{1},\cdots,k_{m})\geq 1-\epsilon&\text{if $s\in Q(T[k_{1},\cdots,k_{m}])$}\\ P_{m}(k_{1},\cdots,k_{m})\leq\epsilon&\text{if $s\notin Q(T[k_{1},\cdots,k_{m}])$}\end{cases}$
Let $\tilde{P}_{m}:=P_{m}-\frac{1}{2}$ and $\epsilon^{\prime}:=\frac{1}{2}-\epsilon$. Interpreting the query $Q$ over $T[k_{1},\cdots,k_{m}]$ implies the following constraints on the sequence of polynomial $\tilde{P}_{m}$:
(3) $\exists\epsilon^{\prime}>0\;\text{such that}\begin{cases}\forall k\in\mathbb{N}^{m},\exists i\in\{1,\cdots,m\},k_{i}=0\implies\tilde{P}_{m}(k)\leq-\epsilon^{\prime}&\\ \forall k\in\mathbb{N}^{m},\forall i\in\{1,\cdots,m\},k_{i}>0\implies\tilde{P}_{m}(k)\geq\epsilon^{\prime}&\end{cases}$
Let $q$ be the degree of $P_{m}$, $q:=\mathsf{deg}(P_{m})=\mathsf{deg}(\tilde{P}_{m})$ and let $S$ be the set of exponents of the monomials of $P_{m}$, i.e.
[…]
First, note that $q\geq 2$ ($\tilde{P}_{m}$ cannot be linear, otherwise its zero locus would contain a union of several hyperplanes as soon as $m\geq 2$). This insures that $\lvert S\rvert\geq 2$.
Then using Lemma 1 (with $p=m$) tells us there exists $\alpha^{*}\in S$ and $u=(u_{1},\cdots,u_{m})\in\mathbb{N}^{m}$ such that for any $\alpha^{\prime}\in S-\{\alpha^{*}\}$,
$\langle\alpha^{*},u\rangle>\langle\alpha^{\prime},u\rangle$
Claim 1: The (univariate) monomial $t^{\langle\alpha^{*},u\rangle}$ is the monomial of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m}})$ of largest degree.
Proof.
$\tilde{P}_{m}=\sum_{\alpha\in S}\gamma_{\alpha}k_{1}^{\alpha_{1}}\cdots k_{m}^{\alpha_{m}}\implies\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},t^{u_{m}})=\sum_{\alpha\in S}\gamma_{\alpha}t^{\langle\alpha,u\rangle}$
Hence, the monomial of largest degree of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m}})$ is the one such that $\langle\alpha,u\rangle$ is (strictly) maximized when $\alpha\in S$. By construction, it is $\alpha^{*}$. ∎
Now, suppose that $\tilde{P}_{m}$ has bounded degree (i.e. there is a uniform bound on the degree of the polynomials $\tilde{P}_{m}$ that is independent of $m$). Then, any monomial of largest degree of $\tilde{P}_{m}$ does not contain all the variables. Without loss of generality (by symmetry of $\tilde{P}_{m}$), we can suppose that the monomial $k_{1}^{\alpha^{*}_{1}}\cdots k_{m}^{\alpha^{*}_{m}}$ does not contain $k_{m}$, i.e. $\alpha^{*}_{m}=0$.
Claim 2: In these conditions, the (univariate) monomial $t^{\langle\alpha^{*},u\rangle}$ is also the monomial of of largest degree of $P_{m}(t^{u_{1}},\allowbreak\cdots,t^{u_{m-1}},0)$.
Proof.
Evaluating $\tilde{P}_{m}$ in $(t^{u_{1}},\cdots,t^{u_{m-1}},0)$ removes the contribution of each monomial of $P_{m}$ containing the last variable, and keeps only the contribution of the monomials containing it:
$\tilde{P}_{m}=\sum_{\alpha\in S}\gamma_{\alpha}k_{1}^{\alpha_{1}}\cdots k_{m}^{\alpha_{m}}\implies\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},0)=\sum_{\begin{subarray}{c}\alpha\in S\alpha=(\alpha_{1},\cdots,\alpha_{m-1},0)\end{subarray}}\gamma_{\alpha}t^{\langle\alpha,u\rangle}$
Therefore, the monomial of largest degree of $\tilde{P}_{m}(t^{u_{1}},\cdots,\allowbreak t^{u_{m-1}},0)$ is the one such that $\langle\alpha,u\rangle$ is (strictly) maximized when $\alpha\in S$ and $\alpha_{m}=0$. By construction, such $\alpha$ is $\alpha^{*}$. ∎
However, Conditions 3 imply that the leading monomial of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m}})$ must have a positive coefficient but the leading monomial of $\tilde{P}_{m}(t^{u_{1}},\cdots,t^{u_{m-1}},0)$ must have a negative coefficient, by taking both limits when $t$ tends to +$\infty$+ ∞ (this limit can be constant). This is a contradiction because Claims 1 and 2 imply that these coefficients are both $\gamma_{\alpha^{*}}$.
Hence if $P_{m}$ expresses $Q$, the leading monomial of $\tilde{P}_{m}$ contains all the variables, and $\mathsf{deg}(P_{m})=\mathsf{deg}(\tilde{P}_{m})\geq m$. This gives a contradiction with Lemma 1 since $P_{m}=\tilde{P}_{m}+\frac{1}{2}$ is supposed to be computed by a GNN. Therefore, no GNN with polynomial activations and aggregations can uniformly express the query $Q$.
6. Discussion and open problems
The expressive capabilities of GNNs are accurately characterized by the color refinement (or Wesfeiler-Leman) algorithm and fragments of the 2-variable counting logic. It is valuable to comprehend the expressivity of machine learning architectures, and in particular those of GNNs, as it can help in selecting an appropriate architecture for a given problem and facilitates the comparison of various architectures and approaches. It is also essential to note that expressivity is just one facet of practical use of GNNs and the related machine learning algorithms. This paper does not delve into other crucial aspects such as the GNN’s ability to generalize from provided data, and the computational efficiency of learning and inference. In particular, we have not investigated the ability of a GNN to learn a logical query from examples (without knowing the query in advance), for instance, from a sample complexity standpoint. We believe that theoretical investigations on the expressivity of GNNs and logical expressivity can also suggest potential avenues to integrate logic-based and statistical reasoning in machine learning methods, and in particular in GNN architectures. These investigations also include questions of independent mathematical interest, some of which remain open. We list some below that are closely related to the results presented in this article:
Question 1: Can GNNs with sigmoidal activations and $\mathsf{agg}=\mathsf{sum}$ can uniformly express GC2 queries?
Question 2: Can GNNs with sigmoidal activations and polynomial activations uniformly express GC2 queries?
We conjecture that the answer to both 1 and 2 is negative. The answer to Question 3 seems less clear.
Question 3: Can GNNs with polynomial activation functions and a non polynomial aggregation function express uniformly GC2 queries?
More generally, we believe to be of interest to characterize the logics expressible by GNNs with certain activations and aggregation functions:
Question 4: What is the fragment of GC2 uniformly expressible by GNNs with polynomial activations and aggregation functions?
Remark 5.
Note that the proof of Theorem 2 can easily be extended to a larger family of queries. Namely, for any integer $p\geq 2$, let
$Q_{p}(s):=\lnot\left(\exists^{\geq 1}x(E(s,x)\land\exists^{\leq(p-1)}sE(x,s))\right)=\forall xE(s,x)\exists^{\geq p}sE(x,s)$
$Q_{p}$ queries if vertex $s$ has neighbors whose degree are all at least $p$. Then any $(Q_{p})_{p\in\mathbb{N}}$ cannot be expressed by any GNN with polynomial activations.
Remark 6.
The proof of Theorem 4 was initially attempted using queries of the form:
$\tilde{Q}_{p}(s)=\lnot\left(\exists^{\geq 1}x(E(s,x)\land\exists^{\geq(p+1)}sE(x,s))\right)=\forall xE(s,x)\exists^{\leq p}sE(x,s)$
Which expresses that all the neighbors of $s$ have degree at most $p$. Note the similarity between $Q_{p}$ from Remark 5 and $\tilde{Q}_{p}$. Although $\tilde{Q}_{p}$ also seems a good candidate that cannot be expressed uniformly by a GNN with polynomial activations and aggregations, we could not conclude with the same approach as in the proof of Theorem 4, due to the following interesting fact:
There exists $\epsilon>0$ and a sequence of symmetric polynomial $(p_{m})_{m\in\mathbb{N}}\in\mathbb{R}[x_{1},\cdots,x_{m}]$ of bounded degree (i.e. there exists an integer $q$ such that for any $m$, $\mathsf{deg}(p_{m})\leq q$) and for any $m,p_{m}$ is greater than $\epsilon$ on the vertices of the unit hypercube $\{0,1\}^{m}$, and less than $-\epsilon$ on all the other points of $\mathbb{N}^{m}$. $p_{m}=1-\sum_{i=1}^{m}x_{i}^{2}+\sum_{i=1}^{m}x_{i}^{4}$ is an example of such sequence of symmetric polynomials.
7. Acknowledgements
The author thanks Amitabh Basu, Edinah K. Gnang, Martin Grohe, Pablo Barceló and Floris Geerts for insightful discussions. The author gratefully acknowledges support from Air Force Office of Scientific Research (AFOSR) grant FA95502010341.
Appendix A Logic background: general definitions
Definition 7.
A first order logic is given by a countable set of symbols, called the alphabet of the logic:
  1. (1)
    Boolean connectives: ¬,,,$\neg$,$\vee$,$\wedge$,$\rightarrow$,$\leftrightarrow$ , ∨ , ∧ , → , ↔
  2. (2)
    Quantifiers: $\forall$,$\exists$ , ∃
  3. (3)
    Equivalence/equality symbol: $\equiv$
  4. (4)
    Variables: $x_{0},x_{1},\ldots$ (finite or countably infinite set)
  5. (5)
    Punctuation: ((((, )))) and ,.
  6. (6)
    1. (a)
      A (possibly empty) set of constant symbols.
    2. (b)
      For every natural number $n\geq 1$, a (possibly empty) set of $n$-ary function symbols.
    3. (c)
      For every natural number $n\geq 1$, a (possibly empty) set of $n$-ary relation symbols.
Remark 7.
Items 1-5 are common to any first order logic. Item 6 changes from one system of logic to another. Example: In Graph theory, the first order logic has:
  • no constant symbols
  • no function symbol
  • a single $2$-ary relation symbol $E$ (which is interpreted as the edge relation between vertices). When graphs are supposed labeled with $\ell$ colors: $\ell$ function symbols $\mathsf{col}_{1},\cdots,\mathsf{col}_{\ell}$. $\mathsf{col}_{i}(v\in G)$ returns $1$ if the color of $v$ is the $i$-th one.
The set of symbols from Item 6 is called the vocabulary of the logic. It will be denoted by $S$ and the first order logic based on $S$ will be denoted by $FO(S)$.
Definition 8.
The set of terms in a given first order logic $FO(S)$ is a set of strings over the alphabet defined inductively as follows:
  1. (1)
    Every variable and constant symbol is a term.
  2. (2)
    If $f$ is an $n$-ary function symbol, and $t_{1},\ldots,t_{n}$ are terms, then $f(t_{1},\ldots,t_{n})$ is a term.
Definition 9.
The set of formulas in a given first order logic is a set of strings over the alphabet defined inductively as follows:
  1. (1)
    If $t_{1},t_{2}$ are terms, then $t_{1}\equiv t_{2}$ is a formula.
  2. (2)
    If $R$ is an $n$-ary relation symbol, and $t_{1},\ldots,t_{n}$ are terms, then $R(t_{1},\ldots,t_{n})$ is a formula.
  3. (3)
    If $\phi$ is a formula, then $\neg\phi$ is a formula.
  4. (4)
    If $\phi_{1},\phi_{2}$ are formulas, then $(\phi_{1}\vee\phi_{2})$, $\phi_{1}\wedge\phi_{2}$, $\phi_{1}\rightarrow\phi_{2}$ and $\phi_{1}\leftrightarrow\phi_{2}$ are formulas.
  5. (5)
    If $\phi$ is a formula and $x$ is a variable, then $\forall x\phi$ and $\exists x\phi$ are formulas.
The set of all variable symbols that appear in a term $t$ will be denoted by $\operatorname{var}(t)$. The set of free variables in a formula is defined using the inductive nature of formulas:
  1. (1)
    $\operatorname{free}(t_{1}\equiv t_{2})=\operatorname{var}(t_{1})\cup\operatorname{var}(t_{2})$
  2. (2)
    $\operatorname{free}(R(t_{1},\ldots,t_{n}))=\operatorname{var}(t_{1})\cup\ldots\cup\operatorname{var}(t_{n})$
  3. (3)
    $\operatorname{free}(\neg\phi)=\operatorname{free}(\phi)$
  4. (4)
[…]
  5. (5)
    $\operatorname{free}(\forall x\phi)=\operatorname{free}(\phi)\setminus\{x\}$
  6. (6)
    $\operatorname{free}(\exists x\phi)=\operatorname{free}(\phi)\setminus\{x\}$
Remark 8.
The same variable symbol may be a free symbol in $\phi$, but appear bound to a quantifier in a subformula of $\phi$.
Definition 10.
The set of sentences in a first order logic are all the formulas with no free variables, i.e., $\{\phi:\operatorname{free}(\phi)=\emptyset\}$.
Definition 11.
Given a first order logic $FO(S)$, an $S$-structure is a pair $\mathcal{U}=(A,I)$ where $A$ is a nonempty set, called the domain/universe of the structure, and $I$ is a map defined on $S$ such that
  1. (1)
    $I(c)$ is an element of $A$ for every constant symbol $c$.
  2. (2)
    $I(f)$ is a function from $A^{n}$ to $A$ for every $n$-ary function symbol $f$.
  3. (3)
    $I(R)$ is a function from $A^{n}$ to $\{0,1\}$ (or equivalently, a subset of $A^{n}$) for every $n$-ary relation symbol $R$.
Given an $S$-structure $\mathcal{U}=(A,I)$ for $FO(S)$, an assignment is a map from the set of variables in the logic to the domain $A$. An interpretation of $FO(S)$ is a pair $(\mathcal{U},\beta)$, where $\mathcal{U}$ is an $S$-structure and $\beta$ is an assignment.
We say that an interpretation $(\mathcal{U},\beta)$ satisfies a formula $\phi$, if this assignment restricted to the free variables in $\phi$ evaluates to 1, using the standard Boolean interpretations of the symbols of the first order logic in Items 1-5 of Definition 7.
Definition 12.
[…]
This is equivalent to the depth of the tree representing the formula, based on the inductive definition. The length/size of the formula is the total number nodes in this tree. Up to constants, this is the number of leaves in the tree, which are called the atoms of the formula.